Question 2: Firms Consider a firm that produces output Y from capital K and labour N using the production iechoolopy Y KNdThe fs capital endowcnt is piven as K 50 Labour is hired to maximize profits. At a wage rate w, the firms labour costs are wN The firms profit (as a function of N is therefore 1. Find the firms labour demand function by maximizing profits and solving the fist order condition for the wage rate w. 2. Plot the labour demand curve (Nd on the horizontal axis, w on the vertical axis). 3. What wage rate would the firm be willing to pay if it were to hire Nd-8 units of labour? (Use your labour demand function!) What would the firms profit be in this case? (Use the profit function!) 4.
Answer #1

a)

Given

\pi(N^{^{d}}) =50^{{1/2}}(N^{^{d}})^{1/2}-wN^{d}

Differentiate with respect to Nd , we get

\partial \pi(N^{^{d}})/\partial N^{d}=50^{1/2}*1/2*(N^{d})^{-1/2}-w

Put  \partial \pi(N^{^{d}})/\partial N^{d}=0 for profit maximization

50^{1/2}*1/2*(N^{d})^{-1/2}-w=0

50^{1/2}*1/2-(N^{d})^{1/2}*w=0

50^{1/2}*1/2=(N^{d})^{1/2}*w

Squaring both sides

\frac{50}{4}=N^{d}*w^{2}

N^{d}=\frac{25}{2w^{2}}

Labor demand function is given by

N^{d}=\frac{25}{2w^{2}}

b)

w Nd=12.5/w2
0.5 50.00
1 12.50
1.5 5.56
2 3.13
2.5 2.00
3 1.39
3.5 1.02
4 0.78
4.5 0.62
5 0.50
5.5 0.41
6 0.35

8Pb2cyF7dbQXMAAAAASUVORK5CYII=

c)

Given Nd =8, We have derived in part(b) that

N^{d}=\frac{25}{2w^{2}}

8=\frac{25}{2w^{2}}

16w^{2}=25

w=\frac{5}{4}=1.25

d)

Firm's Profit in this case is given by

\pi(N^{^{d}}) =50^{{1/2}}(N^{^{d}})^{1/2}-wN^{d}

\pi(N^{^{d}}) =50^{{1/2}}(8)^{1/2}-1.25*8=20-10=10